You want to achieve an oxygen concentration of 80% by blending air and O2 together. To achieve this, you would have an air:O2 ratio of _______.
Answer #1

- To achieve an oxygen of 80%

Use the following formula:

Air litres/ oxygen litres= 100- O2%/ O2% - 21

= 100-80/ 80-21

= 20/59

= 0.3 = 0.3:1

So, air to oxygen ratio is 0.3:1